Let E and F be events for which P(E) = .5, P(F)= .4, and P(E \cap F) = .2

a) are E and F mutually exclusive or independent? (justify mathematically)

b) Find P(E\cup F)

c) Find P(F')

d) Find P(F l E)

e) Find P(E' \cap F)
Answer #1

A) For a mutually exclusive event P(A and B) =0, but here P(E and F)= 0.2. So,can not be a exclusive events.

For a independent events P( A and B)= P(A) * P(B) and P(E and F) =P(E)*P(F)= 0.5*0.4=0.2. So, both are independent events.

B) P(A or B)= P(A) +P(B)- P(A and B)

P(E or F) = P(E) +P(F)- P(E and F) = 0.5+0.4-0.2= 0.7

C) P(F')

P(A') =1-P(A)

P(F')= 1-P(F) =1- 0.4= 0.6

D) P( F | E)

P(E)0.5

For an independent event: P(F|E)= rac{P(E cap F)}{P(E)}=P(F)

E) P(E' cap F)

P(A' cap B)= P(B)-P(Acap B)

P(E' cap F)= P(F)-P(Ecap F)= 0.4-0.2=0.2